You are the supervisor for the production of H 2 S gas in a chemical plant. Your team can use two

possible processes at STP to produce 20 liters of the gas. Your aim is to cut expenses when procuring

reactants. Which of the two processes will your team use and why?


1) FeS(s) + HCl(aq) = FeCl 2 (aq) + H 2 S(g)



1
Expert's answer
2021-02-22T05:48:38-0500

Using the first method, we need x liters of HCl to produce it. The total equation:

FeS(s) + 2HCl(aq) = FeCl2 (aq) + H2S(g)

n(H2S) = 20/22.4 = 0.8928 (moles)


n(HCl) = 2 * 0.8928 = 1.7857 (moles)

m(HCl) = n(HCl) * M(HCl) = 1.7857*36.46 = 65.1 (grams)


M(FeS) = n(FeS) * M(FeS) = 0.8928 * 89.91 = 80.27 (grams)


To sum up, for preparing 20 liters of H2S we need 80 grams of FeS and 65 grams of HCl, which is quite appropriate in terms of expenses. The second reason for the first method is that it allows to get the purified H2S.